linear algebra
prove that 2 to the power of 3(whole number) equals 8.

Wpi 0
Answer #1
Value of (2)^3 = 2 * 2 * 2 = 8

Value of (2)^3 = Multiplying 2 three times = 2 * 2 * 2 = 8